# Optical properties and electronic correlations in LaNiO bilayer nickelates under high pressure

### v2

@@ II.3 Evolution of the optical spectrum under high pressure
Based on these insights, we now predict the pressure dependence of the optical conductivity of LaNiO consistently at  eV and thereby track the changes in the electronic structure across the structural phase transition.
@@
Figure 3(a) shows , , and for the orthorhombic geometry at ambient pressure. In sharp contrast to the pronounced in-plane versus out-of-plane anisotropy, the in-plane anisotropy is rather small. The inset shows that the optical spectrum exhibits a distinct âhumpâ of moderate anisotropy between and  eV. We assign it predominantly to transitions from O (below  eV) to La (above  eV) [Fig. 1(f)]. Since the corresponding TDMs are relatively low, its considerable magnitude is a cumulative effect related to a high density of states.
@@
At 30 GPa [Fig. 3(b)], the in-plane anisotropy vanishes due to the orthorhombic-to-tetragonal phase transition. The in-plane peaks B (), C (), and D ( eV) as well as the out-of-plane peaks A (), B (), and C ( eV) are shifted to higher photon energies. This can be understood from a pressure-enhanced energy difference between the lower and upper Ni states that is accompanied by an increased energy of the Ni states, specifically around the  point [31]. The in-plane peak A vanishes for finite pressure (not visible) due to the quenched octahedral rotations. Simultaneously, the Drude peak is considerably intensified due to the metallization of the Ni states: The DFT plasma frequencies increase from to  eV and from to  eV (see Supplemental Information). The hump experiences a substantial broadening and now extends towards higher energies owing to the pressure-induced increase of the O band width (lowering of the energy onset of the O states) in the valence band [31].
@@
At 50 GPa [Fig. 3(c)], the in-plane peaks B (), C (), and D ( eV) as well as the out-of-plane peaks A (), B (), and C ( eV) are shifted to even higher energies. Concomitantly, the out-of-plane Drude weight is further increased ( eV).
@@
Interestingly, we observe a pressure-induced intensity reduction of the out-of-plane peak B, which directly reflects the partial depletion of the âbondingâ Ni states in the valence band [Fig. 1(f)]. Together with the Drude peak enhancement, this allows us to experimentally track the proposed emergence of a hole pocket [1, 11, 12, 13, 14, 28, 23]), which plays a key role in the suggested superconductivity mechanism [12, 13, 28, 22, 23]. Therefore, the results promote optical spectroscopy as powerful technique to track transitions of the Fermi surface topology in high-pressure experiments.
@@
II.4 Role of the layer stacking and comparison with related Ruddlesden-Popper nickelates
@@
Finally, we discuss the results for the bilayer compound LaNiO [Fig. 4(c)] in the broader context of related Ruddlesden-Popper (RP, NiO) and reduced Ruddlesden-Popper (RRP, NiO) nickelates.
@@
We observe comparable shapes of the optical conductivity for the three RP members [Fig. 4(a-c)] shown. The trilayer nickelate LaNiO [, Fig. 4(a)] has recently been proposed to be superconducting [60, 61]. The characteristic out-of-plane peak, which is found to universally stem from transitions between Ni--derived quantum-well states [cf. Fig. 1(f)], is lowered to  eV due to the reduced quantum confinement relative to the bilayer compound.
@@
An interesting variant of the bilayer geometry is the recently observed 1-3 stacking [42, 43] [Fig. 4(b)]. We find it to be metastable and  meVf.u. above the 2-2 ground state. The simulated spectra strongly resemble their analogs. In particular, the characteristic out-of-plane peak appears at the same energy as in LaNiO, yet with reduced intensity. Unique manifestations of the single-layer Ni states in the spectrum are not observed.
@@
We find the best agreement of the experimental spectrum with our theoretical results for LaNiO in the bilayer 2-2 stacking [Fig. 4(c)]. In contrast, in the predicted spectrum for the 1-3 stacking, the in-plane peak B is significantly lower than the experimental data and peak D is not visible at all, whereas peak C and the spectral weight below  eV are overestimated. We speculate that 2-2 and 1-3 phases may coexist in experimental samples. Intriguingly, the results establish as ideal observable to quantify their relative volume fraction due to the highly characteristic peak energies.
@@
Complete reduction of the apical oxygen in LaNiO results in the RRP structure LaNiO [62, 63] [Fig. 4(d)]. The formal Ni valence jumps from to due to the released electrons, which we find to be largely accommodated by the Ni states. Hence, the spectral weight up to  eV is strongly suppressed; specifically, the characteristic peaks in and vanish. A double peak appears in between - eV, which is far less intense than in the RP nickelates and of different origin; we attribute it to transitions between Ni and La states.
@@
The end member of the RRP series is the infinite-layer geometry [Fig. 4(e)], in which the Ni states are fully occupied [51, 64]. Here we consider NdNiO [Fig. 4(e)], which hosts superconductivity in film geometry [4]. The in-plane optical conductivity agrees reasonably with earlier DFTDMFT simulations for LaNiO [59] (inset), particular in the interband regime, which corroborates the validity of our approach. In addition, we predict the out-of-plane component, which uncovers a further reduced anisotropy. With respect to LaNiO, the interband onset of shifts to lower energies and now coincides with at  eV. While the characteristic peaks are absent, we find the higher-energy spectral weight to be enhanced in the RRP versus RP compounds.
@@
III Summary
The optical properties of LaNiO bilayer nickelates were investigated by performing density functional theory simulations including a Coulomb repulsion term. We analyzed specifically the microscopic origin of the interband transitions, which reflect the relative energies of the active Ni orbitals and their distance to the O states in the valence band, and uncovered a surprisingly strong anisotropy of the optical response that reverses with frequency. The optical spectrum was found to be considerably impacted by on-site correlation effects, and very good agreement with experiment at ambient pressure is attained for  eV. This value simultaneously provides Ni energies consistent with recent angle-resolved photoemission spectroscopy as well as lattice parameters that are in close agreement with x-ray diffraction.
@@ III Summary
Furthermore, we discussed the topical question of the 2-2 versus 1-3 layer stacking in LaNiO. The out-of-plane optical conductivity emerges as ideal observable to quantify the relative volume fraction of these possibly coexisting phases due to their highly characteristic peaks.
Finally, we broadened the perspective by comparing the bilayer nickelate to LaNiO, LaNiO, and NdNiO, which unveiled that each material exhibits its unique optical signature. We identified fundamental differences between Ruddlesden-Popper and reduced Ruddlesden-Popper nickelates, but also general trends that relate to the formal Ni valence.
In conclusion, the present comprehensive study establishes the notion of rather moderate electronic correlations in LaNiO. The uncovered trends and their microscopic analysis may guide future experiments, support the quest towards understanding the origin of superconductivity in this high- nickelate, and accelerate the discovery of related compounds with enhanced properties.
Acknowledgements.
We thank Prof. Roser ValentĂ for helpful discussions. This work was supported by the National Science Foundation, Grant No. NSF-DMR-2118718. L. F. acknowledges support by the European Unionâs Horizon 2020 research and innovation programme through the Marie SkĹodowska-Curie grant SuperCoop (Grant No. 838526).COMPETING INTERESTS
The authors declare no competing interests.
Contributions
@@ Acknowledgements.
Methods
We performed first-principles simulations in the framework of density functional theory (DFT [65]) as implemented in the Vienna Ab initio Simulation Package (VASP) [66, 67], employing a wave-function cutoff of  eV. Exchange and correlations were described by using the generalized gradient approximation as parameterized by Perdew, Burke, and Ernzerhof [68]. Static correlation effects were considered within the DFT formalism [69, 70], using  eV at the Ni sites unless stated otherwise.
To account for octahedral tilts and oxygen vacancies, the LaNiO bilayer nickelates () were modeled by using orthorhombic 48-atom unit cells. The Brillouin zone was sampled employing Monkhorst-Pack [71] and -centered -point grids in conjunction with a Gaussian smearing of  mRy. The lattice parameters , and as well as the internal ionic positions were accurately optimized in DFT under zero and finite external pressure, reducing ionic forces below  mRy/a.u. We proceeded analogously for LaNiO, LaNiO, and NdNiO.
where models the oxygen-rich limit 111The well-known overbinding of gas-phase O molecules in DFT necessitates a correction of , which we performed such as to reproduce the experimental O binding energy of  eV [47, 55, 74, 49, 75, 57]..
Here, label the different spatial directions, Â denotes the unit cell volume, and and enumerate the conduction and valence band states, respectively. These interband contributions are complemented by the Drude peak:
The plasma frequency squared can be derived directly from the DFT electronic structure (see Supplemental Information). In contrast, the lifetime is considerably more difficult to obtain. We employ  meV here, motivated by recent experimental Drude peak analysis for LaNiO [30] and DFTDMFT results for LaNiO [59]. Subsequently, the real part of the optical conductivity is obtained via

### v3

@@ II.3 Evolution of the optical spectrum under high pressure
Based on these insights, we now predict the pressure dependence of the optical conductivity of La3Ni2O7 consistently at eV and thereby track the changes in the electronic structure across the structural phase transition.
@@
Figure 3(a) shows , , and for the orthorhombic geometry at ambient pressure. In sharp contrast to the pronounced in-plane versus out-of-plane anisotropy, the in-plane anisotropy is rather small. The inset shows that the optical spectrum exhibits a distinct ’hump’ of moderate anisotropy between and eV. We assign it predominantly to transitions from O (below eV) to La (above eV) [Fig. 1(f)]. Since the corresponding TDMs are relatively low, its considerable magnitude is a cumulative effect related to a high density of states.
@@
At 30 GPa [Fig. 3(b)], the in-plane anisotropy vanishes due to the orthorhombic-to-tetragonal phase transition. The in-plane peaks B (), C (), and D ( eV) as well as the out-of-plane peaks A (), B (), and C ( eV) are shifted to higher photon energies. This can be understood from a pressure-enhanced energy difference between the lower and upper Ni states that is accompanied by an increased energy of the Ni states, specifically around the point [31]. The in-plane peak A vanishes for finite pressure (not visible) due to the quenched octahedral rotations. Simultaneously, the Drude peak is considerably intensified due to the metallization of the Ni states [ pocket, Fig. 3(d)]: The DFT plasma frequencies increase from to eV2 and from to eV2 (see Supplementary Information). The hump experiences a substantial broadening and now extends towards higher energies owing to the pressure-induced increase of the O band width (lowering of the energy onset of the O states) in the valence band [31].
@@
At 50 GPa [Fig. 3(c)], the in-plane peaks B (), C (), and D ( eV) as well as the out-of-plane peaks A (), B (), and C ( eV) are shifted to even higher energies. Concomitantly, the out-of-plane Drude weight is further increased ( eV2).
@@
Interestingly, we observe a pressure-induced intensity reduction of the out-of-plane peak B, which directly reflects the partial depletion of the ’bonding’ Ni states in the valence band [Figs. 3(e) and 1(f)]. Simultaneously, a new transition from eV to these depleted states becomes active [Fig. 3(e)]. Together with the Drude peak enhancement, this allows us to experimentally track the proposed emergence of a hole pocket [1, 11, 12, 13, 14, 28, 23]), which plays a key role in the suggested superconductivity mechanism [12, 13, 28, 22, 23]. Therefore, the results promote optical spectroscopy as powerful technique to track transitions of the Fermi surface topology in high-pressure experiments.
II.4 Optical signature of different magnetic phases
Recent measurements suggested the presence of a spin-density wave (SDW) in La3Ni2O7 at ambient pressure [65, 38]. However, nuclear diffraction and nuclear magnetic resonance studies found no long-range magnetic order [66, 67]. Motivated by this topical discussion, we explore the optical spectra for different magnetic phases in Fig. 4, which have been obtained by performing simulations for large supercells. We find a unique optical signature with strong in-plane versus out-of-plane anisotropy for each magnetic phase. For instance, the prominent peak in (A, B) is shifted to higher energies (- eV) for the antiferromagnetic (AFM) stripe order suggested in Ref. [23] and the AFM- order (with and without site disproportionation, i.e., an in-plane checkerboard charge and spin modulation [31]) and is less pronounced than in the nonmagnetic case. For all AFM phases [Fig. 4(c)-(f)], the Drude peak is negligible or fully quenched, and the interband onset of is found at higher energies than for .
The bi-stripe AFM order superimposed by Ni2+/Ni3+ charge order [Fig. 4(f)] has been proposed recently by experiment [68] and theory [69] as good representation of the SDW in La3Ni2O7. Similar to AFM- order with site disproportionation, the Drude peak is absent due to the emergence of a finite band gap.
Interestingly, the experimental curve, which has been measured at K [30] above the SDW transition, agrees best with the result of our nonmagnetic simulation, judging from the overall shape and peak positions of the interband spectrum together with the presence of a Drude peak. Notably, the sample remains metallic down to low temperatures [30]. These results will be helpful in the interpretation of future low-temperature studies.
@@ II.4 Optical signature of different magnetic phases
II.5 Role of the layer stacking and comparison with related Ruddlesden-Popper nickelates
@@
Finally, we discuss the results for the bilayer compound La3Ni2O7 [Fig. 5(c)] in the broader context of related Ruddlesden-Popper (RP, NinO3n+1) and reduced Ruddlesden-Popper (RRP, NinO2n+2) nickelates.
@@
We observe comparable shapes of the optical conductivity for the three RP members [Fig. 5(a-c)] shown. The trilayer nickelate La4Ni3O10 [, Fig. 5(a)] has recently been proposed to be superconducting [71, 72]. The characteristic out-of-plane peak, which is found to universally stem from transitions between Ni--derived quantum-well states [cf. Fig. 1(f)], is lowered to eV due to the reduced quantum confinement relative to the bilayer compound.
@@
An interesting variant of the bilayer geometry is the recently observed 1-3 stacking [42, 43] [Fig. 5(b)]. We find it to be metastable and meVf.u. above the 2-2 ground state. The simulated spectra strongly resemble their analogs. In particular, the characteristic out-of-plane peak appears at the same energy as in La4Ni3O10, yet with reduced intensity. Unique manifestations of the single-layer Ni states in the spectrum are not observed.
@@
We find the best agreement of the experimental spectrum with our theoretical results for La3Ni2O7 in the bilayer 2-2 stacking [Fig. 5(c)]. In contrast, in the predicted spectrum for the 1-3 stacking, the in-plane peak B is significantly lower than the experimental data and peak D is not visible at all, whereas peak C and the spectral weight below eV are overestimated. We speculate that 2-2 and 1-3 phases may coexist in experimental samples. Intriguingly, the results establish as ideal observable to quantify their relative volume fraction due to the highly characteristic peak energies.
@@
Complete reduction of the apical oxygen in La3Ni2O7 results in the RRP structure La3Ni2O6 [73, 74] [Fig. 5(d)]. The formal Ni valence jumps from to due to the released electrons, which we find to be largely accommodated by the Ni states. Hence, the spectral weight up to eV is strongly suppressed; specifically, the characteristic peaks in and vanish. A double peak appears in between - eV, which is far less intense than in the RP nickelates and of different origin; we attribute it to transitions between Ni and La states.
@@
The end member of the RRP series is the infinite-layer geometry [Fig. 5(e)], in which the Ni states are fully occupied [53, 75]. Here we consider NdNiO2 [Fig. 5(e)], which hosts superconductivity in film geometry [4]. The in-plane optical conductivity agrees reasonably with earlier DFTDMFT simulations for LaNiO2 [70] (inset), particular in the interband regime, which corroborates the validity of our approach. In addition, we predict the out-of-plane component, which uncovers a further reduced anisotropy. With respect to La3Ni2O6, the interband onset of shifts to lower energies and now coincides with at eV. While the characteristic peaks are absent, we find the higher-energy spectral weight to be enhanced in the RRP versus RP compounds.
@@
III Discussion
The optical properties of La3Ni2O7 bilayer nickelates were investigated by performing density functional theory simulations including a Coulomb repulsion term. We analyzed specifically the microscopic origin of the interband transitions, which reflect the relative energies of the active Ni orbitals and their distance to the O states in the valence band, and uncovered a surprisingly strong anisotropy of the optical response that reverses with frequency. The optical spectrum was found to be considerably impacted by on-site correlation effects, and very good agreement with experiment at ambient pressure is attained for eV. This value simultaneously provides Ni energies consistent with recent angle-resolved photoemission spectroscopy as well as lattice parameters that are in close agreement with x-ray diffraction.
@@ III Discussion
Moreover, we identified the unique optical signature of different magnetic phases in La3Ni2O7 at ambient pressure, including very recent models of spin-density waves. These insights will be helpful in the interpretation of future low-temperature studies.
Furthermore, we discussed the topical question of the 2-2 versus 1-3 layer stacking in La3Ni2O7. The out-of-plane optical conductivity emerges as ideal observable to quantify the relative volume fraction of these possibly coexisting phases due to their highly characteristic peaks.
Finally, we broadened the perspective by comparing the bilayer nickelate to La4Ni3O10, La3Ni2O6, and NdNiO2, which unveiled that each material exhibits its unique optical signature. We identified fundamental differences between Ruddlesden-Popper and reduced Ruddlesden-Popper nickelates, but also general trends that relate to the formal Ni valence.
In conclusion, the present comprehensive study establishes the notion of rather moderate electronic correlations in La3Ni2O7. The uncovered trends and their microscopic analysis may guide future experiments, support the quest towards understanding the origin of superconductivity in this high- nickelate, and accelerate the discovery of related compounds with enhanced properties.
Methods
Density functional theory calculations
We performed first-principles simulations in the framework of (spin-polarized) density functional theory (DFT [76]) as implemented in the Vienna Ab initio Simulation Package (VASP) [77, 78], employing a wave-function cutoff of eV. Exchange and correlations were described by using the generalized gradient approximation as parameterized by Perdew, Burke, and Ernzerhof [79]. Static correlation effects were considered within the DFT formalism [80, 81], using eV at the Ni sites unless stated otherwise.
To account for octahedral tilts and oxygen vacancies, the La3Ni2O7-δ bilayer nickelates () were modeled by using orthorhombic 48-atom unit cells. The Brillouin zone was sampled employing Monkhorst-Pack [82] and -centered -point grids in conjunction with a Gaussian smearing of mRy. The lattice parameters , and as well as the internal ionic positions were accurately optimized in DFT under zero and finite external pressure, reducing ionic forces below mRy/a.u. We proceeded analogously for La4Ni3O10, La3Ni2O6, and NdNiO2.
where models the oxygen-rich limit. The well-known overbinding of gas-phase O2 molecules in DFT necessitates a correction of , which we performed such as to reproduce the experimental O2 binding energy of eV [48, 83, 84, 50, 85, 86].
Here, label the different spatial directions, denotes the unit cell volume, and and enumerate the conduction and valence band states, respectively. These interband contributions are complemented by the Drude peak:
We predict the plasma frequency squared directly from the DFT electronic structure (see Supplementary Information). In contrast, the lifetime is considerably more difficult to obtain. We employ meV here, motivated by recent experimental Drude peak analysis for La3Ni2O7 [30] and DFTDMFT results for LaNiO2 [70]. Subsequently, the real part of the optical conductivity is obtained via
@@ Methods
Acknowledgements.
We thank Prof. Roser Valentí for helpful discussions. This work was supported by the National Science Foundation, Grant No. NSF-DMR-2118718. L. F. acknowledges support by the European Union’s Horizon 2020 research and innovation programme through the Marie Skłodowska-Curie grant SuperCoop (Grant No. 838526).Author Contributions
Competing Interests
The authors declare no competing interests.